Hi Can you assist with below and show all the work?

16.   Calculating Monthly Mortgage Payments: Based on Exhibit 7-7, what would be the monthly mortgage payments for each of the following situations?

a.         A $100,000, 15-year loan at 7.5% APR compounded semi-annually
b.         A $200,000, 25-year loan at 9% APR compounded semi-annually
c.         A $165,000, 20-year loan at 10% APR compounded semi-annually

What relationship exists between the length of the loan and the monthly payment? How does the mortgage rate affect the monthly payment?
Answer #1

А 1 amount time 3 period 4 rate 5 rate pmt 100000 15 years 30 semiannually 7.50% APR 3.75% semiannually $5,608.76 9 amount 20

100000 15 1 amount 2 time 3 period 4 rate 5 rate 6 pmt =+B2*2 0.075 =+B4/2 =PMT(B5,B3,-B1,,0) years semiannually APR semiannu